OpenStudy (anonymous):

Evaluate the intergral from 0 to 3 of (x^3 -6x)

OpenStudy (earthcitizen):

\[\int\limits_{0}^{3}x ^{3}-6x dx\] =\[ (3)^4/4-3 (3)^2 - 0\]

OpenStudy (earthcitizen):

\[\lim_{3 \rightarrow 0} x^4/4-3 x^2\]

OpenStudy (earthcitizen):

\[\lim_{0 \rightarrow 3} x^4/4-3 x^2\]=\[(3)^4/4-3 (3)^2 - 0\] = -(-27/4) = 27/4

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty} {(18/n^4)[(n(n+1))/2]^{2} - 54/n ^{2} [n(n+1)/2}\]